Alyeska Services Company, a division of a major oil company, provides various services to the operators of the North Slope oil field in Alaska. Data concerning the most recent year appear below: Sales $ 18,800,000 Net operating income $ 5,300,000 Average operating assets $ 36,500,000 Required: 1. Compute the margin for Alyeska Services Company. (Round your answer to 2 decimal places.) 2. Compute the turnover for Alyeska Services Company. (Round your answer to 2 decimal places.) 3. Compute the return on investment (ROI) for Alyeska Services Company. (Round your intermediate calculations and final answer to 2 decimal places.)

1.Margin%

2.Turnover

3.ROI%
Answer #1

Margin = Net operating income/Sales

= 5,300,000/18,800,000

= 28.19%

.

Turnover = Sales/Average operating assets

= 18,800,000/36,500,000

= 0.52

.

Roi = Margin * Turnover

= 28.19*0.52

= 14.66%

Note : I have rounded some numbers. Comment if you face any issues